# Watson-Crick Conjugates of Words and Languages

### v3

@@ 3. Some Important Properties of set of all $\theta$-conjugates of a Word
Authors in [29] proved that if $C_{\theta}(u)=C_{\theta}(v)$ and $|C_{\theta}(u)|\geq 3$ for $u,v\in\Sigma^{*}$, then either $u=v$ or $u=v^{R}$ (Proposition 3.15 in [29]). However, the detailed structures of $u$ and $v$ were provided within the proof of Proposition 3.15 in [29]. Since we will later use these structural details, we have explicitly included them in Theorem 3.10. Additionally, to ensure clarity and verifiability, we have added the proof of these structural details in the appendix.
@@
For $u,v\in\Sigma^{k}$, $u\neq v$ and $|C_{\theta}(u)|=|C_{\theta}(v)|=2$, consider $C_{\theta}(u)=C_{\theta}(v)$. Then by Theorem 3.10, $v=\theta(u)$ where $\theta(u)\neq u$.
@@
Consider $u$ and $v$ are two distinct words of same length such that $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 3$. Then, $C_{\theta}(u)=C_{\theta}(v)$ if and only if $u=a^{m+1}ba^{m}$ and $v=a^{m}ba^{m+1}$ where $m\geq 1$, $a,b\in\Sigma$, $a\neq b$, $\theta(a)=a$ and $\theta(b)=b$.
@@
For $u,v\in\Sigma^{*}$, $u\neq v$ and $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 3$, consider $C_{\theta}(u)=C_{\theta}(v)$. Then, by Theorem 3.10,
@@
where $a\in\Sigma$, $n_{1},n_{2}\geq 1$, $n_{1}\neq n_{2}$, $u^{\prime}\in\Sigma^{+}$, $a\notin Pref(u^{\prime})$, $a\notin Suf(u^{\prime})$, $u^{\prime}=u^{\prime R}$ and $\theta(c)=c~{}\forall c\in\operatorname{Alph}(u)$.
@@
Conversely, let $u=a^{m+1}ba^{m}$ and $v=a^{m}ba^{m+1}$ where $m\geq 1$, $a,b\in\Sigma$, $\theta(a)=a$, $\theta(b)=b$ and $a\neq b$. Then $C_{\theta}(u)=\{a^{m+1+j}ba^{m-j}~{}:~{}0\leq j\leq m\}\cup\{v\}$, and $C_{\theta}(v)=\{v\}\cup\{a^{m+i}ba^{m+1-i}~{}:~{}1\leq i\leq m+1\}\cup\{a^{m+1}ba^{m}\}$. Since $a^{m+1}ba^{m}\in\{a^{m+i}ba^{m+1-i}~{}:~{}1\leq i\leq m+1\}$, we have $C_{\theta}(v)=\{v\}\cup\{a^{m+i}ba^{m+1-i}~{}:~{}1\leq i\leq m+1\}$. Consider $i=i^{\prime}+1$. Then, $11\leq i\leq m+$ 1 implies $0\leq i^{\prime}\leq m$. Thus, $C_{\theta}(v)=\{v\}\cup\{a^{m+i^{\prime}+1}ba^{m-i^{\prime}}~{}:~{}0\leq i^{\prime}\leq m\}$. Therefore, $C_{\theta}(u)=C_{\theta}(v)$. ∎
It is well known that for given $u,v\in\Sigma^{*}$, $C(u)=C(v)$ if and only if $u\in C(v)$. But this statement is not true in general for $\theta$-conjugates of a word. Propositions 3.13 and 3.14 describe necessary and sufficient conditions for the equality of $C_{\theta}(u)$ and $C_{\theta}(v)$ for distinct words $u$ and $v$ with $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 2$. Also, it is clear from Propositions 3.13 and 3.14 that if sets of all $\theta$-conjugates of two distinct words $u\in\Sigma^{\geq 2}$ and $v\in\Sigma^{\geq 2}$ are equal, then $u,v$ must be binary words and $\theta(c)=c$ for all $c\in\operatorname{Alph}(u)\cup\operatorname{Alph}(v)$. Now, combining Propositions 3.13 and 3.14, we have the following characterizations of $u$ and $v$ such that sets of all $\theta$-conjugates of $u$ and $v$ are equal.
@@
4. $\theta$-conjugates of a Language And Closure Properties of $\theta$-conjugate Operation
Conjugate operation or cyclic shift operation ([15]) is a unary operation on formal languages defined as $C(L)=\bigcup\limits_{w\in L}C(w)$ for some $L\subseteq\Sigma^{*}$. Here, we call $C(L)$ as the set of all conjugates of the language $L$. In [15], the authors showed that the family of regular languages is closed under the conjugate operation. In addition, the author of [30] demonstrated that if $L$ is a context-free language, then $C(L)$ is also context-free. Families of context-sensitive and recursively enumerable languages have also been shown to be closed under the conjugate operation by the author of [6]. Now, the definition of $\theta$-conjugate of a word can be naturally extended to a language. In this section, we first define the $\theta$-conjugate operation on languages and investigate the closure property of certain families of languages under the $\theta$-conjugate operation.
@@ 4. $\theta$-conjugates of a Language And Closure Properties of $\theta$-conjugate Operation
First, we show that the family of regular languages is closed under $\theta$-conjugate operation. The proof is very similar to the argument presented by the authors of [15] on the closure of regular languages under the conjugate operation.
@@
Consider a deterministic finite automaton (DFA) $A$ which accepts $L$. Let $Q^{\prime}$, $q_{0}$, and $F$ be the set of all states, the initial state, and the set of all final states of $A$, respectively. Now, for each state $q\in Q^{\prime}$, we first construct two DFAs $C_{q}$ and $B_{q}$ with same states and same transition functions as $A$ where $q_{0}$ is the initial state of $C_{q}$, $\{q\}$ is the set of all final states of $C_{q}$, $q$ is the initial state of $B_{q}$ and $F$ is the set of all final states of $B_{q}$. Consider a morphic involution $\psi:\Sigma^{*}\rightarrow\Sigma^{*}$ such that $\psi(a)=\theta(a)$ for all $a\in\Sigma$. Then, as the family of regular languages is closed under reversal and morphism [27], there exists a DFA $B_{q}^{\prime}$ which accepts $\psi((L(B_{q}))^{R})$, i.e., $\theta(L(B_{q}))$ (by Lemma 4.2). Now, by definition, a string $w$ is in $C_{\theta}(L)$ if and only if $w=\theta(v)u$ for $uv\in L$. Then, in the accepting computation of $A$ on $uv$, there exists an intermediate state $q$ of $A$ such that the computation of $A$ on $u$ ends in the state $q$, while $v$ is accepted by $A$ starting from $q$. Then, $\theta(v)$ is accepted by $B_{q}^{\prime}$ and $u$ is accepted by $C_{q}$ and $C_{\theta}(L)=\bigcup_{q\in Q^{\prime}}L(B_{q}^{\prime})L(C_{q})$. Since the family of regular languages is closed under concatenation and union [27], $C_{\theta}(L)$ is regular. ∎
@@
[35] If $L$ is context-free, then there exists a constant $n$ such that for all $z\in L$ with $|z|\geq n$, there exists a decomposition $z=uvwxy$ with $|vwx|\leq n$ and $|vx|\geq 1$ such that for all $i\geq 0$, we have $uv^{i}wx^{i}y\in L$.
@@
Consider the following context-free language $L=\{a^{n}b^{k}c^{k}d^{n}\;|\;n,k\geq 1\}$ over $\Sigma=\{a,b,c,d\}$. Suppose $\theta(e)=e$ for all $e\in\Sigma$. Then, $C_{\theta}(L)=\{d^{i}a^{n}b^{k}c^{k}d^{n-i}\;|\;n,k\geq 1,0\leq i\leq n\}$cup $\{d^{n}c^{j}a^{n}b^{k}c^{k-j}\;|\;n,k\geq 1,0<j\leq k\}$cup $\{d^{n}c^{k}b^{l}a^{n}b^{k-l}\;|\;n,k\geq 1,0<l\leq k\}$. Consider the regular language $L^{\prime}=\{d^{n_{1}}c^{m_{1}}a^{m_{2}}b^{n_{2}}\;|\;n_{1},m_{1},\\ m_{2},n_{2}\geq 1\}$. Now, if $C_{\theta}(L)$ is context-free, then as the intersection of a context-free language and a regular language is context-free [35], $C_{\theta}(L)\cap L^{\prime}=\{d^{n}c^{k}a^{n}b^{k}~{}|~{}n,k\geq 1\}$ is context-free. But by pumping lemma for context-free languages (Theorem 4.4), $\{d^{n}c^{k}a^{n}b^{k}~{}|~{}n,k\geq 1\}$ is not context-free (proof is in Appendix), which is a contradiction. Thus, $C_{\theta}(L)$ is not context-free. ∎
@@
Let $L_{1}\subseteq\{a,b\}^{+}$ be a context-sensitive language. Since the family of context-sensitive languages is closed under reversal [33], $L_{1}^{R}$ is also context-sensitive. For $\Sigma=\{a,b\}$, let $\psi:\Sigma^{*}\rightarrow\Sigma^{*}$ be a morphic involution such that $\psi(c)=\theta(c)$ for all $c\in\Sigma$. Consider
@@
Now, as $L_{1}^{R}$ is a context-sensitive language, there exists a linear bounded automaton $M=(Q^{\prime},\Sigma\cup\{<,>\},\Gamma,\delta,q^{\prime}_{0},<,>,B,F)$ that accepts $L_{1}^{R}$ ([26]) and uses $(k+2)\cdot|z|+2$ tape cells for any $z\in\Sigma^{+}$ with the initial configuration $q_{0}^{\prime}<\underset{((|z|-1)\text{ no. of blank cells})}{BB\cdots B}z\underset{((k\cdot|z|+1)\text{ no. of blank cells})}{BB\cdots B}>$ where $k$ is a constant.
Using $M$, we now construct a linear bounded automaton $M^{\prime}$ which accepts $L^{\prime}$. Let $M^{\prime}=(Q^{\prime\prime},\Sigma\cup\{<,>\},\Gamma^{\prime},\delta^{\prime},q_{0},<,>,B,F)$ where $Q^{\prime\prime}=Q^{\prime}\cup\{q_{i}~{}|~{}0\leq i\leq 15\}$, $\Gamma^{\prime}=\Gamma\cup\{\#,X\}$, $\delta^{\prime}(q,c)=\delta(q,c)$ for all $q\in Q^{\prime}$, $c\in\Gamma$. For $M^{\prime}$, we will use $(k+2)\cdot|w|+2$ tape cells for each $w\in\Sigma^{*}$. Also, for any $w\in\Sigma^{*}$, we consider the initial configuration of $M^{\prime}$ as $q_{0}<\underset{(|w|\text{ no. of blank cells})}{BB\cdots B}w\underset{(k\cdot|w|\text{ no. of blank cells})}{BB\cdots B}>$. In the transition diagram (Figure 4), we have discussed the transition rules of $M^{\prime}$ in details. Here we have assumed that $q^{\prime}_{0}=q_{15}$. Also, in Figure 4, the label $a_{1}|b_{1},R$ (resp. $L$ or $N$) on the arc from state $q_{i}$ to state $q_{j}$ signifies that when the machine is in state $q_{i}$ with the head reading $a_{1}$, the machine goes to state $q_{j}$, writes $b_{1}$, and moves the head to the right (resp. left or stay in its current position).
Suppose $\alpha\in\Sigma^{+}$ is an input to $M^{\prime}$. Then, for every factorization $xy$ of $\alpha$ where $x,y\in\Sigma^{*}$, the linear bounded automaton $M^{\prime}$ computes the following: $q_{0}<\underset{(|\alpha|\text{ no. of blank cells})}{BB\cdots B}xy\underset{(k\cdot|\alpha|\text{ no. of blank cells})}{BB\cdots B}>{\vdash}^{*}_{M^{\prime}}~{}q_{15}<\underset{(|\alpha|-1\text{ no. of blank cells})}{BB\cdots B}\psi(x)y^{R}\underset{(k\cdot|\alpha|+1\text{ no. of blank cells})}{BB\cdots B}>$, which is the initial configuration of $M$ for $\psi(x)y^{R}$.
Figure 4. Transition diagram for $M^{\prime}$
@@
5. Iterated $\theta$-conjugate of Words And Languages
@@
$|u|_{a,\theta(a)}=\begin{cases}|u|_{a}+|u|_{\theta(a)}&\text{if }\theta(a)\neq a\\ |u|_{a}&\text{if }\theta(a)=a.\end{cases}$
@@
Since the conjugacy relation is an equivalence relation [27], $u\in C(w)$ implies $C(u)=C(w)$. Then, $C(C(w))=C^{2}(w)=C(w)$, $C(C^{2}(w))=C^{3}(w)=C(w),\ldots,C(C^{i-1}(w))=C^{i}(w)=C(w)$ for all $i\geq 1$. But this is not true for $\theta$-conjugates of a word. For an example, $ba\in{C_{\theta}}^{2}(ab)$ but $ba\notin C_{\theta}(ab)$ where $\theta(a)=b$. We now show that for a given $w\in\Sigma^{n}$, there exists an integer $m$ depending on $n$ such that ${C_{\theta}}^{m}(w)$ contains all words $u$ such that $|u|_{a,\theta(a)}$ is equal to $|w|_{a,\theta(a)}$ for all $a\in\Sigma$. To prove this, we need the following results.
@@
From the definition of iterated $\theta$-conjugate of $w$, it is clear that ${C_{\theta}}^{4n-6}(w)\subseteq{C_{\theta}}^{*}(w)$. Consider $x\in{C_{\theta}}^{*}(w)$. Then, for some integer $k$, $x\in{C_{\theta}}^{k}(w)$. If $k<4n-6$, then by Lemma 5.2, ${C_{\theta}}^{k}(w)\subseteq{C_{\theta}}^{4n-6}(w)$. This implies $x\in{C_{\theta}}^{4n-6}(w)$. If $k\geq 4n-6$, then by Proposition 5.6, ${C_{\theta}}^{4n-6}(w)={C_{\theta}}^{k}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}\;\forall a\in\Sigma\}$. This implies $x\in{C_{\theta}}^{4n-6}(w)$. Thus, ${C_{\theta}}^{*}(w)\subseteq{C_{\theta}}^{4n-6}(w)$. Hence, ${C_{\theta}}^{4n-6}(w)={C_{\theta}}^{*}(w)$. ∎
We now provide an example to show that for a word $w$ of length $n\geq 3$, the bound $64n-$ 6 is optimal. Additionally, we present an example to show that the smallest $i$ such that ${C_{\theta}}^{i}(w)={C_{\theta}}^{*}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}\;\forall a\in\Sigma\}$ may or may not be equal to $64n-$ 6.
@@
    Hence, ${C_{\theta}}^{*}(w)={C_{\theta}}^{6}(w)$. Note that, for $m=6=4|w|-6$, ${C_{\theta}}^{*}(w)={C_{\theta}}^{m}(w)$.
  2. (2)
@@
    Hence, ${C_{\theta}}^{*}(w)={C_{\theta}}^{4}(w)$. Note that, for $m=4<4|u|-6=6$, ${C_{\theta}}^{*}(w)={C_{\theta}}^{m}(w)$.
@@
For $w\in\Sigma^{n}$, let $\operatorname{Alph}(w)=A_{1}\cup A_{2}\cup B$ where $A_{1}\cap A_{2}=\emptyset$, $A_{1}\cap B=\emptyset$, $A_{2}\cap B=\emptyset$, $B=\{b:\theta(b)=b\}$, and $A_{2}=\{\theta(a):a\in A_{1}\text{ and }\theta(a)\in\operatorname{Alph}(w)\}$. Consider $A_{1}=\{a_{1},a_{2},\cdots,a_{l}\}$ and $B=\{b_{1},b_{2},\cdots,b_{r}\}$ for some non-negative integers $l$ and $r$. For each $a\in A_{1}$ and $b\in B$, let $|w|_{a,\theta(a)}=n_{a}$ and $|w|_{b,\theta(b)}=m_{b}$, respectively. Consider $\sum\limits_{a\in A_{1}}n_{a}=n^{\prime}$. Then
@@
Consider $\operatorname{Alph}(w)=A_{1}\cup A_{2}\cup B$ where $A_{1}\cap A_{2}=\emptyset$, $A_{1}\cap B=\emptyset$, $A_{2}\cap B=\emptyset$, $B=\{b:\theta(b)=b\}=\{b_{1},b_{2},\cdots,b_{r}\}$, $A_{1}=\{a_{1},a_{2},\cdots,a_{l}\}$ and $A_{2}=\{\theta(a):a\in A_{1}\text{ and }\theta(a)\in\operatorname{Alph}(w)\}$. Then $A_{1}$ is the collection of all those elements of $\operatorname{Alph}(w)$ such that for any two distinct indices $i$ and $j$ with $1\leq i,j\leq l$, we have $a_{i}\neq\theta(a_{i})$, $a_{j}\neq\theta(a_{j})$ and $a_{i}\neq\theta(a_{j})$.

### v5

@@ 3 Some Important Properties of set of all $\theta$-conjugates of a Word
Mahalingam et al. (2022) proved that if $C_{\theta}(u)=C_{\theta}(v)$ and $|C_{\theta}(u)|\geq 3$ for $u,v\in\Sigma^{*}$, then either $u=v$ or $u=v^{R}$ (Proposition 3.15 in Mahalingam et al. (2022)). However, the detailed structures of $u$ and $v$ were provided within the proof of Proposition 3.15 in Mahalingam et al. (2022). Since we will later use these structural details, we have explicitly included them in Theorem 3.10. Additionally, to ensure clarity and verifiability, we have added the proof of these structural details in the appendix.
@@
For $\in\Sigma^{k}, uvu\neq v$ and $|C_{\theta}(u)|=|C_{\theta}(v)|=2$, consider $C_{\theta}(u)=C_{\theta}(v)$. Then by Theorem 3.10, $v=\theta(u)$ where $\theta(u)\neq u$.
@@
Consider $u$ and $v$ are two distinct words of same length such that $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 3$. Then, $C_{\theta}(u)=C_{\theta}(v)$ if and only if $u=a^{m+1}ba^{m}$ and $v=a^{m}ba^{m+1}$ where $m\geq 1$, $\in\Sigma, aba\neq b$, $\theta(a)=a$ and $\theta(b)=b$.
@@
For $\in\Sigma^{*}, uvu\neq v$ and $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 3$, consider $C_{\theta}(u)=C_{\theta}(v)$. Then, by Theorem 3.10,
@@
where $a\in\Sigma$, $n_{1},n_{2}\geq 1$, $n_{1}\neq n_{2}$, $u^{\prime}\in\Sigma^{+}$, $a\notin Pref(u^{\prime})$, $a\notin Suf(u^{\prime})$, $u^{\prime}=u^{\prime R}$ and $\theta(c)=c~\forall c\in\textit{Alph}(u)$.
@@
Conversely, let $u=a^{m+1}ba^{m}$ and $v=a^{m}ba^{m+1}$ where $m\geq 1$, $a,b\in\Sigma$, $\theta(a)=a$, $\theta(b)=b$ and $a\neq b$. Then $C_{\theta}(u)=\{a^{m+1+j}ba^{m-j}~:~0\leq j\leq m\}\cup\{v\}$, and $C_{\theta}(v)=\{v\}\cup\{a^{m+i}ba^{m+1-i}~:~1\leq i\leq m+1\}\cup\{a^{m+1}ba^{m}\}$. Since $a^{m+1}ba^{m}\in\{a^{m+i}ba^{m+1-i}~:~1\leq i\leq m+1\}$, we have $C_{\theta}(v)=\{v\}\cup\{a^{m+i}ba^{m+1-i}~:~1\leq i\leq m+1\}$. Consider $i=i^{\prime}+1$. Then, $1\leq i\leq m+1$ implies $0\leq i^{\prime}\leq m$. Thus, $C_{\theta}(v)=\{v\}\cup\{a^{m+i^{\prime}+1}ba^{m-i^{\prime}}~:~0\leq i^{\prime}\leq m\}$. Therefore, $C_{\theta}(u)=C_{\theta}(v)$. ∎
It is well known that for given $u,v\in\Sigma^{*}$, $C(u)=C(v)$ if and only if $u\in C(v)$. But this statement is not true in general for $\theta$-conjugates of a word. Propositions 3.13 and 3.14 describe necessary and sufficient conditions for the equality of $C_{\theta}(u)$ and $C_{\theta}(v)$ for distinct words $u$ and $v$ with $|C_{\theta}(u)|=|C_{\theta}(v)|\geq 2$. Also, it is clear from Propositions 3.13 and 3.14 that if sets of all $\theta$-conjugates of two distinct words $u\in\Sigma^{\geq 2}$ and $v\in\Sigma^{\geq 2}$ are equal, then $u,v$ must be binary words and $\theta(c)=c$ for all $c\in\textit{Alph}(u)\cup\textit{Alph}(v)$. Now, combining Propositions 3.13 and 3.14, we have the following characterizations of $u$ and $v$ such that sets of all $\theta$-conjugates of $u$ and $v$ are equal.
@@
4 $\theta$-conjugates of a Language And Closure Properties of
$\theta$-conjugate Operation
Conjugate operation or cyclic shift operation (Jirásková and Okhotin (2008)) is a unary operation on formal languages defined as $C(L)=\bigcup\limits_{w\in L}C(w)$ for some $L\subseteq\Sigma^{*}$. Here, we call $C(L)$ as the set of all conjugates of the language $L$. Jirásková and Okhotin (2008) showed that the family of regular languages is closed under the conjugate operation. In addition, Oshiba (1972) demonstrated that if $L$ is a context-free language, then $C(L)$ is also context-free. Families of context-sensitive and recursively enumerable languages have also been shown to be closed under the conjugate operation by Brandstädt (1981). Now, the definition of $\theta$-conjugate of a word can be naturally extended to a language. In this section, we first define the $\theta$-conjugate operation on languages and investigate the closure property of certain families of languages under the $\theta$-conjugate operation.
@@ 4 $\theta$-conjugates of a Language And Closure Properties of
First, we show that the family of regular languages is closed under $\theta$-conjugate operation. The proof is very similar to the argument presented by Jirásková and Okhotin (2008) on the closure of regular languages under the conjugate operation.
@@
Consider a deterministic finite automaton (DFA) $A$ which accepts $L$. Let $Q^{\prime}$, $q_{0}$, and $F$ be the set of all states, the initial state, and the set of all final states of $A$, respectively. Now, for each state $q\in Q^{\prime}$, we first construct two DFAs $C_{q}$ and $B_{q}$ with same states and same transition functions as $A$ where $q_{0}$ is the initial state of $C_{q}$, $\{q\}$ is the set of all final states of $C_{q}$, $q$ is the initial state of $B_{q}$ and $F$ is the set of all final states of $B_{q}$. Consider a morphic involution $\psi:\Sigma^{*}\rightarrow\Sigma^{*}$ such that $\psi(a)=\theta(a)$ for all $a\in\Sigma$. Then, as the family of regular languages is closed under reversal and morphism Lothaire (1997), there exists a DFA $B_{q}^{\prime}$ which accepts $\psi((L(B_{q}))^{R})$, i.e., $\theta(L(B_{q}))$ (by Lemma 4.2). Now, by definition, a string $w$ is in $C_{\theta}(L)$ if and only if $w=\theta(v)u$ for $uv\in L$. Then, in the accepting computation of $A$ on $uv$, there exists an intermediate state $q$ of $A$ such that the computation of $A$ on $u$ ends in the state $q$, while $v$ is accepted by $A$ starting from $q$. Then, $\theta(v)$ is accepted by $B_{q}^{\prime}$ and $u$ is accepted by $C_{q}$ and $C_{\theta}(L)=\bigcup_{q\in Q^{\prime}}L(B_{q}^{\prime})L(C_{q})$. Since the family of regular languages is closed under concatenation and union Lothaire (1997), $C_{\theta}(L)$ is regular. ∎
@@
Shallit (2008) If $L$ is context-free, then there exists a constant $n$ such that for all $z\in L$ with $|z|\geq n$, there exists a decomposition $z=uvwxy$ with $|vwx|\leq n$ and $|vx|\geq 1$ such that for all $i\geq 0$, we have $uv^{i}wx^{i}y\in L$.
@@
Consider the following context-free language $L=\{a^{n}b^{k}c^{k}d^{n}\;|\;n,k\geq 1\}$ over $\Sigma=\{a,b,c,d\}$. Suppose $\theta(e)=e$ for all $e\in\Sigma$. Then, $C_{\theta}(L)=\{d^{i}a^{n}b^{k}c^{k}d^{n-i}\;|\;n,k\geq 1,0\leq i\leq n\}$ $\cup$ $\{d^{n}c^{j}a^{n}b^{k}c^{k-j}\;|\;n,k\geq 1,0<j\leq k\}$ $\cup$ $\{d^{n}c^{k}b^{l}a^{n}b^{k-l}\;|\;n,k\geq 1,0<l\leq k\}$. Consider the regular language $L^{\prime}=\{d^{n_{1}}c^{m_{1}}a^{m_{2}}b^{n_{2}}\;|\;n_{1},m_{1},\\ m_{2},n_{2}\geq 1\}$. Now, if $C_{\theta}(L)$ is context-free, then as the intersection of a context-free language and a regular language is context-free Shallit (2008), $C_{\theta}(L)\cap L^{\prime}=\{d^{n}c^{k}a^{n}b^{k}~|~n,k\geq 1\}$ is context-free. But by pumping lemma for context-free languages (Theorem 4.4), $\{d^{n}c^{k}a^{n}b^{k}~|~n,k\geq 1\}$ is not context-free (proof is in Appendix), which is a contradiction. Thus, $C_{\theta}(L)$ is not context-free. ∎
@@
Let $L_{1}\subseteq\{a,b\}^{+}$ be a context-sensitive language. Since the family of context-sensitive languages is closed under reversal Rozenberg and Salomaa (Eds.), $L_{1}^{R}$ is also context-sensitive. For $\Sigma=\{a,b\}$, let $\psi:\Sigma^{*}\rightarrow\Sigma^{*}$ be a morphic involution such that $\psi(c)=\theta(c)$ for all $c\in\Sigma$. Consider
@@
[…]
Figure 4: Transition diagram for $M^{\prime}$
@@
5 Iterated $\theta$-conjugate of Words And Languages
@@
[…]
@@
Since the conjugacy relation is an equivalence relation Lothaire (1997), $u\in C(w)$ implies $C(u)=C(w)$. Then, $C(C(w))=C^{2}(w)=C(w)$, $C(C^{2}(w))=C^{3}(w)=C(w),\ldots,C(C^{i-1}(w))=C^{i}(w)=C(w)$ for all $i\geq 1$. But this is not true for $\theta$-conjugates of a word. For an example, $ba\in{C_{\theta}}^{2}(ab)$ but $ba\notin C_{\theta}(ab)$ where $\theta(a)=b$. We now show that for a given $w\in\Sigma^{n}$, there exists an integer $m$ depending on $n$ such that ${C_{\theta}}^{m}(w)$ contains all words $u$ such that $|u|_{a,\theta(a)}$ is equal to $|w|_{a,\theta(a)}$ for all $a\in\Sigma$. To prove this, we need the following results.
@@
From the definition of iterated $\theta$-conjugate of $w$, it is clear that ${C_{\theta}}^{4n-6}(w)\subseteq{C_{\theta}}^{*}(w)$. Consider $x\in{C_{\theta}}^{*}(w)$. Then, for some integer $k$, $x\in{C_{\theta}}^{k}(w)$. If k<4n6k<4n-6, then by Lemma 5.2, ${C_{\theta}}^{k}(w)\subseteq{C_{\theta}}^{4n-6}(w)$. This implies $x\in{C_{\theta}}^{4n-6}(w)$. If $k\geq 4n-6$, then by Proposition 5.6, ${C_{\theta}}^{4n-6}(w)={C_{\theta}}^{k}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}\;\forall a\in\Sigma\}$. This implies $x\in{C_{\theta}}^{4n-6}(w)$. Thus, ${C_{\theta}}^{*}(w)\subseteq{C_{\theta}}^{4n-6}(w)$. Hence, ${C_{\theta}}^{4n-6}(w)={C_{\theta}}^{*}(w)$. ∎
We now provide an example to show that for a word $w$ of length $n\geq 3$, the bound 4n64n-6 is optimal. Additionally, we present an example to show that the smallest $i$ such that ${C_{\theta}}^{i}(w)={C_{\theta}}^{*}(w)=\{u\;|\;|u|_{a,\theta(a)}=|w|_{a,\theta(a)}\;\forall a\in\Sigma\}$ may or may not be equal to 4n64n-6.
@@
    Hence, ${C_{\theta}}^{*}(w)={C_{\theta}}^{6}(w)$. Note that, for m=6=4|w|6m=6=4|w|-6, ${C_{\theta}}^{*}(w)={C_{\theta}}^{m}(w)$.
  2. 2.
@@
    Hence, ${C_{\theta}}^{*}(w)={C_{\theta}}^{4}(w)$. Note that, for m=4<4|u|6=6m=4<4|u|-6=6, ${C_{\theta}}^{*}(w)={C_{\theta}}^{m}(w)$.
@@
For $w\in\Sigma^{n}$, let $\textit{Alph}(w)=A_{1}\cup A_{2}\cup B$ where $A_{1}\cap A_{2}=\emptyset$, $A_{1}\cap B=\emptyset$, $A_{2}\cap B=\emptyset$, $B=\{b:\theta(b)=b\}$, and $A_{2}=\{\theta(a):a\in A_{1}\text{ and }\theta(a)\in\textit{Alph}(w)\}$. Consider $A_{1}=\{a_{1},a_{2},\cdots,a_{l}\}$ and $B=\{b_{1},b_{2},\cdots,b_{r}\}$ for some non-negative integers $l$ and $r$. For each $a\in A_{1}$ and $b\in B$, let $|w|_{a,\theta(a)}=n_{a}$ and $|w|_{b,\theta(b)}=m_{b}$, respectively. Consider $\sum\limits_{a\in A_{1}}n_{a}=n^{\prime}$. Then
@@
Consider $\textit{Alph}(w)=A_{1}\cup A_{2}\cup B$ where $A_{1}\cap A_{2}=\emptyset$, $A_{1}\cap B=\emptyset$, $A_{2}\cap B=\emptyset$, $B=\{b:\theta(b)=b\}=\{b_{1},b_{2},\cdots,b_{r}\}$, $A_{1}=\{a_{1},a_{2},\cdots,a_{l}\}$ and $A_{2}=\{\theta(a):a\in A_{1}\text{ and }\theta(a)\in\textit{Alph}(w)\}$. Then $A_{1}$ is the collection of all those elements of $\textit{Alph}(w)$ such that for any two distinct indices $i$ and $j$ with $1\leq i,j\leq l$, we have $a_{i}\neq\theta(a_{i})$, $a_{j}\neq\theta(a_{j})$ and $a_{i}\neq\theta(a_{j})$.